# Gravity as a deformed topological gauge theory

### v6

@@ 3.2.2 The total action
In particular, this shows that by imposing specific constraints on the components $\alpha$ and $\lambda$ of the Cartan connection $\varpi$ for $G=SO(4,2)/\{\pm I\}$, one can recover the action for the subgroups $SO(4,1),~SO(3,2),~\text{or }ISO(3,1)$ already considered in Section 2. Indeed, one has with the respective constraints:
$\{\lambda=0\text{ and }\alpha=\dfrac{k}{2}\beta\}\,\cap\begin{array}[]{ll}\left\{\begin{array}[]{ll}sign(k\ell\ell^{\prime})=-,\quad\Lambda_{0}=-\dfrac{3k}{\ell\ell^{\prime}}>0&\Leftrightarrow G=SO(4,1)\\ k=0&\Leftrightarrow G=ISO(3,1)\\ sign(k\ell\ell^{\prime})=+,\quad\Lambda_{0}=-\dfrac{3k}{\ell\ell^{\prime}}<0&\Leftrightarrow G=SO(3,2).\end{array}\right.\end{array}$ (37)
888-dimensional manifold $\widetilde{\mathscr{M}}$ action
In the case $\alpha\neq\dfrac{k}{2}\beta$ corresponding to the 888-dimensional manifold $\widetilde{\mathscr{M}}$ ($\alpha$ is related to the special conformal transformations) the action (1) becomes:
$S_{G}[\varpi]$ =~$=\int_{\tilde{\mathscr{M}}}\bigl{(}\overbrace{r\det(\mathds{1}+\dfrac{\bar{\Omega}}{2\pi})}^{P(\bar{\Omega})}+y\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})\bigr{)}$ (38)
=~$=\int_{\tilde{\mathscr{M}}}\Bigl{(}-\dfrac{r}{128\pi^{2}}\bigl{(}\operatorname{Tr}(\bar{\Omega}^{2})^{2}-2\operatorname{Tr}(\bar{\Omega}^{4})\bigr{)}-\dfrac{y}{128\pi^{2}}\bigl{(}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{2})^{2}-2\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{4})\bigr{)}\Bigr{)}$
Study of action 38 shall be the subject of another article.
4 Topological total action
@@
[…]
@@
In the middle equation, the first term vanishes thanks to the $\mathfrak{h}$-part of the Bianchi identity, namely, $d\bar{\Omega}_{\mathfrak{h}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{h}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}]=0$, the second term is nothing but $ad_{\varpi_{\mathfrak{h}}}P(\bar{\Omega}_{\mathfrak{h}}^{n})$ which vanishes by $H$-invariance. The third term will be under discussion whether $P$ is $G$-invariant or if solely its $H$-invariance will be considered.
@@
The use of the Bianchi identity for $\bar{\Omega}_{t}$ in the right hand side, relies on the Ad-invariance of $P$, which in our situation at hand, depends on either the $G$ or the $H$-invariance of $P$.
Case 1. For $P\in{\cal I}^{n}(G)$, the $Ad(G)$-invariance tells us that
$ad_{\varpi_{t}}P(\varpi,\bar{\Omega}_{t}^{n-1})=P([\varpi_{t},\varpi],\bar{\Omega}_{t}^{n-1})-(n-1)P(\varpi,[\varpi_{t},\bar{\Omega}_{t}],\bar{\Omega}_{t}^{n-2})=0$ (4)
@@
$n\,dP(\varpi,\bar{\Omega}_{t}^{n-1})$ $=nP(d\varpi,\bar{\Omega}_{t}^{n-1})-n(n-1)P(\varpi,d\bar{\Omega}_{t},\bar{\Omega}_{t}^{n-2})+ad_{\varpi_{t}}P(\varpi,\bar{\Omega}_{t}^{n-1})$
@@
[…]
Restriction of $P\in{\cal I}^{n}(G)$ to the subgroup $H\subset G$, requires to consider the $\mathfrak{h}$-part of the curvature from the decomposition $\bar{\Omega}=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}$ to obtain the polynomial $P(\bar{\Omega}_{\mathfrak{h}}^{n})$. Going back to (4), since $P\in{\cal I}^{n}(G)$, one can use the $ad(\mathfrak{m})$-invariance,
$ad_{\varpi_{\mathfrak{m}}}P(\bar{\Omega}_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}^{n-1})=P([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})+(n-1)P(\bar{\Omega}_{\mathfrak{m}},[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}],\bar{\Omega}_{\mathfrak{h}}^{n-2})=0.$
This latter relation allows to relax some conditions on torsion of the Cartan connection in order to secure (4) to be 0 ($P(\bar{\Omega}_{\mathfrak{h}}^{n})$ is closed) if one wishes to avoid to consider the $[\mathfrak{m},\mathfrak{m}]$ contributions. Indeed, the Bianchi identity $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ infers that $[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ if the covariant derivative of the torsion $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]=0$. One thus gets
@@
[…]
@@
Case 2. For $P\in{\cal I}^{n}(H)$, only the $H$-invariance can be taken into account. Accordingly, one must restrict eqs(2) and (3) to the $\bar{\Omega}_{\mathfrak{h}}$ and $\bar{\Omega}_{t,\mathfrak{h}}$ components only and for which only the $\mathfrak{h}$-part of the Bianchi identity, $(D\bar{\Omega})_{\mathfrak{h}}=0$, can be used. Together with the result (4), all in all, this gives
@@
where the last two terms could be gathered as the result of $n\,ad_{\varpi_{t,m}}P(\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}^{n-1})$, see (4). But since $P\in{\cal I}^{n}(H)$, the invariance of $P$ under the $\mathfrak{m}$-part of the symmetric Lie algebra $\mathfrak{g}$ no longer holds true. ∎
@@
Corollary 4.1.1.
Another way to reach conditions for the Chern-Weil theorem in both cases is to consider the unique (up to isomorphism) [^22] See [11, Lemma 6.4] and [31, Proposition 4.1]. mutation $\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ such that $[\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=0$.
In the Lorentzian case of Section 2, this mutation corresponds to setting either $k_{1}$ or $k_{2}=0$, thus obtaining the Minkowski geometry.
On the other hand, for the Lorentz×$\times$×Weyl geometry embedded into the mutated Möbius geometry which is a symmetric Lie algebra, the mutation map giving rise (up to isomorphism) to an abelian reductive decomposition could be
@@
where $[\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=[\mathfrak{m}^{\prime\prime},\mathfrak{m}^{\prime\prime}]=0$ and $\mathfrak{m}\simeq\mathfrak{m}^{\prime}\simeq\mathfrak{m}^{\prime\prime}$ as $\mathfrak{h}$-modules. They correspond respectively to (4) with $\bar{\gamma}_{1}=\bar{\gamma}_{2}=0$ or with $\gamma_{1}=\gamma_{2}=0$.
@@
$[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}]$ =$=\begin{cases}\dfrac{k}{\ell^{2}}\begin{pmatrix}\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}\\ \\ \dfrac{k}{\ell\ell^{\prime}}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}\end{cases}$
@@
[…]
For example, taking the Lorentzian geometry of section 2 we can build the total action: q
@@
where we tuned $v$ such that $v,v(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})$ tend towards 0. In this way, we retrieve (if we assume $\mathscr{M}$ has no boundary or when neglecting boundary terms) the equations of motion of section 2 with $\Lambda_{0}\rightarrow 0$. Interestingly, we can make $\Lambda_{0}\rightarrow 0^{\pm}$ such that we end up with a dS or AdS spacetime. To some extent, this limit seems to correspond to the interpretation as a "regulator" of the bare cosmological constant in the case of AdS spacetime [20].
@@
One is still free to set $v$ such that the coupling constants of the Euler, Pontrjagin and Nieh-Yan terms in the action are respectively $-v\dfrac{3}{64\pi G\Lambda_{0}}$, $v(\dfrac{r}{16\pi^{2}}+\dfrac{3}{32\pi G\Lambda_{0}\gamma})$, $\dfrac{v\Lambda_{0}r}{12\pi^{2}}$ and have non-vanishing values. While the other terms that perturb topological invariance asymptotically vanish.
5 Conclusion
@@
$\mathdutchcal{P}(\bar{\Omega})$   Nieh-Yan + Pontrjagin of R               +$d\lambda\wedge d\lambda$ Completion of MM gravity to LQG    with topological BI parameter       and dilational kinetic term
@@
The second example deals with the Möbius geometry. In this case restricting the geometry to a 444-dimensional submanifold $\mathscr{M}$ yields the same results as in the previous class of examples if one drops out dilations. Adding to the original action (3.2) an interaction $T^{a}\wedge\lambda\wedge\beta_{a}$ between dilations and torsion instead leads to Einstein’s equations (modified by the Holst term) with an additional source term (due to the dilational part of the algebra) for the curvature depending on a particular variation of spin density $-\dfrac{2\pi G}{3}\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;kc}$ and on torsion (if one does not assume $T=0$) when coupled to matter.

### v7

@@ III.2.2 The total action
In particular, this shows that by imposing specific constraints on the components $\alpha$ and $\lambda$ of the Cartan connection $\varpi$ for $G=SO(4,2)/\{\pm I\}$, one can recover the action for the subgroups $SO(4,1),~SO(3,2),~\text{or }ISO(3,1)$ already considered in Section II. Indeed, one has with the respective constraints:
[…]
88-dimensional manifold $\widetilde{\mathscr{M}}$ action
In the case $\alpha\neq\dfrac{k}{2}\beta$ corresponding to the $8$-dimensional manifold $\widetilde{\mathscr{M}}$ ($\alpha$ is related to the special conformal transformations) the action (65) becomes:
$S_{G}[\varpi]$ $=\int_{\tilde{\mathscr{M}}}\bigl(\overbrace{r\det(\mathds{1}+\dfrac{\bar{\Omega}}{2\pi})}^{P(\bar{\Omega})}+y\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})\bigr)$ (102)
$=\int_{\tilde{\mathscr{M}}}\Bigl(-\dfrac{r}{128\pi^{2}}\bigl(\operatorname{Tr}(\bar{\Omega}^{2})^{2}-2\operatorname{Tr}(\bar{\Omega}^{4})\bigr)-\dfrac{y}{128\pi^{2}}\bigl(\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{2})^{2}-2\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{4})\bigr)\Bigr)$
Study of action 102 shall be the subject of another article.
IV Topological total action
@@
$(b)$ $\ P(\bar{\Omega}_{\mathfrak{h}}^{n})+\sum_{k=1}^{n}\begin{pmatrix}n\\ k\end{pmatrix}P(\bar{\Omega}_{\mathfrak{h}}^{n-k},\bar{\Omega}_{\mathfrak{m}}^{k})=dQ_{2n-1}(\varpi)\ \mathrm{with}\ Q_{2n-1}(\varpi)=n\int_{0}^{1}dtP(\varpi,\bar{\Omega}^{n-1}_{t}).$
@@
In the middle equation, the first term vanishes thanks to the $\mathfrak{h}$-part of the Bianchi identity, namely, $d\bar{\Omega}_{\mathfrak{h}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{h}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}]=0$, the second term is nothing but $\operatorname{ad}_{\varpi_{\mathfrak{h}}}P(\bar{\Omega}_{\mathfrak{h}}^{n})$ which vanishes by $H$-invariance. The third term will be under discussion whether $P$ is $G$-invariant or if solely its $H$-invariance will be considered.
@@
The use of the Bianchi identity for $\bar{\Omega}_{t}$ in the right hand side, relies on the $\operatorname{Ad}$-invariance of $P$, which in our situation at hand, depends on either the $G$ or the $H$-invariance of $P$.
Case 1. For $P\in{\cal I}^{n}(G)$, the $\operatorname{Ad}(G)$-invariance tells us that
$\operatorname{ad}_{\varpi_{t}}P(\varpi,\bar{\Omega}_{t}^{n-1})=P([\varpi_{t},\varpi],\bar{\Omega}_{t}^{n-1})-(n-1)P(\varpi,[\varpi_{t},\bar{\Omega}_{t}],\bar{\Omega}_{t}^{n-2})=0$ (106)
@@
$n\,dP(\varpi,\bar{\Omega}_{t}^{n-1})$ $=nP(d\varpi,\bar{\Omega}_{t}^{n-1})-n(n-1)P(\varpi,d\bar{\Omega}_{t},\bar{\Omega}_{t}^{n-2})+\operatorname{ad}_{\varpi_{t}}P(\varpi,\bar{\Omega}_{t}^{n-1})$
@@
$P(\bar{\Omega}^{n})=\int_{0}^{1}\frac{d}{dt}P(\bar{\Omega}_{t}^{n})dt=\int_{0}^{1}ndP(\varpi,\bar{\Omega}_{t}^{n-1})dt=dQ_{2n-1}(\varpi).$ (107)
Restriction of $P\in{\cal I}^{n}(G)$ to the subgroup $H\subset G$, requires to consider the $\mathfrak{h}$-part of the curvature from the decomposition $\bar{\Omega}=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}$ to obtain the polynomial $P(\bar{\Omega}_{\mathfrak{h}}^{n})$. Going back to (IV), since $P\in{\cal I}^{n}(G)$, one can use the $\operatorname{ad}(\mathfrak{m})$-invariance,
$\operatorname{ad}_{\varpi_{\mathfrak{m}}}P(\bar{\Omega}_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}^{n-1})=P([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})+(n-1)P(\bar{\Omega}_{\mathfrak{m}},[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}],\bar{\Omega}_{\mathfrak{h}}^{n-2})=0.$
This latter relation allows to relax some conditions on torsion of the Cartan connection in order to secure (IV) to be 0 ($P(\bar{\Omega}_{\mathfrak{h}}^{n})$ is closed) if one wishes to avoid to consider the $\mathfrak{m},\mathfrak{m}]$ contributions. Indeed, the Bianchi identity $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ infers that $\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ if the covariant derivative of the torsion $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]=0$. One thus gets
@@
$(ii)$ $\ P(\bar{\Omega}_{\mathfrak{h}}^{n})+\sum_{k=1}^{n}\begin{pmatrix}n\\ k\end{pmatrix}P(\bar{\Omega}_{\mathfrak{h}}^{n-k},\bar{\Omega}_{\mathfrak{m}}^{k})=dQ_{2n-1}(\varpi)\ \mathrm{with}\ Q_{2n-1}(\varpi)=n\int_{0}^{1}dtP(\varpi,\bar{\Omega}^{n-1}_{t})$
@@
Case 2. For $P\in{\cal I}^{n}(H)$, only the $H$-invariance can be taken into account. Accordingly, one must restrict eqs(104) and (105) to the $\bar{\Omega}_{\mathfrak{h}}$ and $\bar{\Omega}_{t,\mathfrak{h}}$ components only and for which only the $\mathfrak{h}$-part of the Bianchi identity, $(D\bar{\Omega})_{\mathfrak{h}}=0$, can be used. Together with the result (IV), all in all, this gives
@@
where the last two terms could be gathered as the result of $n\,\operatorname{ad}_{\varpi_{t,m}}P(\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}^{n-1})$, see (106). But since $P\in{\cal I}^{n}(H)$, the invariance of $P$ under the $\mathfrak{m}$-part of the symmetric Lie algebra $\mathfrak{g}$ no longer holds true. ∎
As can be seen in $(a),(b)(c)$ and $(d)$, the vanishing of the $\mathfrak{m},\mathfrak{m}]$ commutators removes the obstruction to the Chern-Weil theorem in the case of symmetric Lie algebras. Furthermore, as stated in [35, Proposition 4.1] there is an equivalence between a symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ with $\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ and an abelian reductive decomposition $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ with $\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=0$ through a mutation map and where $\mathfrak{m}\simeq\mathfrak{m}^{\prime}$ as $\mathfrak{h}$-modules, (i.e. $\mathfrak{h},\mathfrak{m}]\subset\mathfrak{m}$ and $\mathfrak{h},\mathfrak{m}^{\prime}]\subset\mathfrak{m}^{\prime}$ and $\mathfrak{m}\simeq\mathfrak{m}^{\prime}$ as vector spaces, see also [11, Lemma 6.4, p.220]). Exploiting this fact, one has
Corollary IV.1.1.
Another way to reach conditions for the Chern-Weil theorem in both cases is to consider the unique (up to isomorphism) [^22] See [11, Lemma 6.4] and [35, Proposition 4.1]. mutation $\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ such that $\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=0$.
In the Lorentzian case of Section II, this mutation corresponds to setting either $k_{1}$ or $k_{2}=0$, thus obtaining the Minkowski geometry.
@@
where $\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=[\mathfrak{m}^{\prime\prime},\mathfrak{m}^{\prime\prime}]=0$ and $\mathfrak{m}\simeq\mathfrak{m}^{\prime}\simeq\mathfrak{m}^{\prime\prime}$ as $\mathfrak{h}$-modules. They correspond respectively to (59) with $\bar{\gamma}_{1}=\bar{\gamma}_{2}=0$ or with $\gamma_{1}=\gamma_{2}=0$.
@@
$[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}]$ $=\begin{cases}\dfrac{k}{\ell^{2}}\begin{pmatrix}\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}\\ \\ \dfrac{k}{\ell\ell^{\prime}}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}\end{cases}$
@@
This shows that these $\mathfrak{m},\mathfrak{m}]$ contributions to $\mathfrak{h}$ can be rendered very small in the limit $\Lambda_{0}\to 0$ [^23] Which is a good approximation if one considers $\Lambda_{\text{exp}}\approx\Lambda_{0}$. provided that the gauge fields and their derivatives have compact supports on $\mathscr{M}$. This leads to a perturbative topological gauge theory in the parameter $\Lambda_{0}$ corresponding to a bare cosmological constant. This can also be seen as a physical implementation of the mutations. In this type of deformed topological gauge theories, the small observed value of the cosmological constant therefore finds a novel interpretation as a perturbatively topological action, with perturbative expansion in the cosmological constant.
For example, taking the Lorentzian geometry of section II we can build the total action:
@@
where we tuned $v$ such that $v,v(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})$ tend towards 0. In this way, we retrieve (if we assume $\mathscr{M}$ has no boundary or when neglecting boundary terms) the equations of motion of section II with $\Lambda_{0}\rightarrow 0$. Interestingly, we can make $\Lambda_{0}\rightarrow 0^{\pm}$ such that we end up with a dS or AdS spacetime. To some extent, this limit seems to correspond to the interpretation as a "regulator" of the bare cosmological constant in the case of AdS spacetime [21].
@@
One is still free to set $v$ such that the coupling constants of the Euler, Pontrjagin and Nieh-Yan terms in the action are respectively $v\dfrac{3}{64\pi G\Lambda_{0}}$, $v(\dfrac{r}{16\pi^{2}}+\dfrac{3}{32\pi G\Lambda_{0}\gamma})$, $\dfrac{v\Lambda_{0}r}{12\pi^{2}}$ and have non-vanishing values. While the other terms that perturb topological invariance asymptotically vanish.
V Conclusion
@@
$\mathdutchcal{P}(\bar{\Omega})$   Nieh-Yan + Pontrjagin of R               $d\lambda\wedge d\lambda$ Completion of MM gravity to LQG    with topological BI parameter       and dilational kinetic term
@@
The second example deals with the Möbius geometry. In this case restricting the geometry to a $4$-dimensional submanifold $\mathscr{M}$ yields the same results as in the previous class of examples if one drops out dilations. Adding to the original action (III.2) an interaction $T^{a}\wedge\lambda\wedge\beta_{a}$ between dilations and torsion instead leads to Einstein’s equations (modified by the Holst term) with an additional source term (due to the dilational part of the algebra) for the curvature depending on a particular variation of spin density $\dfrac{\pi G}{3}\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;kc}$ and on torsion (if one does not assume $T=0$) when coupled to matter.
@@
We also showed the construction of the action motivated by topological arguments leads to constrained couplings of the Euler/Gauss-Bonnet and Pontrjagin densities now expressed in terms of the previous four parameters corresponding to $G$ (gravitational coupling), $\gamma$ (BI parameter), $r$ (Nieh-Yan coupling) and $\Lambda_{0}$ (bare cosmological constant). The values of the constrained couplings are consistant with the specific values allowing for finite Noether charges as described in [21, 20].
In the future we think studying the actions described in this manuscript in terms of BV equivalence [26] could also give very interesting results. The formulation of the deformed topological action (8) in terms of deformations of Chern characteristic classes also motivates us to investigate it as a possible generalization of the Kodama state in gravitational theories.